# Frankenstein Optimizer: Harnessing the Potential by Revisiting Optimization Tricks

### v1

@@ Abstract
Gradient-based optimization drives the unprecedented performance of modern deep neural network models across diverse applications. Adaptive algorithms have accelerated neural network training due to their rapid convergence rates; however, they struggle to find “flat minima” reliably, resulting in suboptimal generalization compared to stochastic gradient descent (SGD). By revisiting various adaptive algorithms’ mechanisms, we propose the Frankenstein optimizer, which combines their advantages. The proposed Frankenstein dynamically adjusts first- and second-momentum coefficients according to the optimizer’s current state to directly maintain consistent learning dynamics and immediately reflect sudden gradient changes. Extensive experiments across several research domains such as computer vision, natural language processing, few-shot learning, and scientific simulations show that Frankenstein surpasses existing adaptive algorithms and SGD empirically regarding convergence speed and generalization performance. Furthermore, this research deepens our understanding of adaptive algorithms through centered kernel alignment analysis and loss landscape visualization during the learning process.
@@ 2.2 Second momentum $v_{t}$ update with dynamic EMA
Starting with the basic RMSProp strategy, which applies a simple Exponential Moving Average (EMA) to the squared gradients, several adaptations have been developed (Appendix B). These methods depends on a fixed $\beta_{2}$ parameter throughout the training process. In contrast, this work proposes a strategy that also relies on squared gradients as its foundation but applies an EMA coefficient that adapts based on both current and historical states. This coefficient is determined by whether the $P$ factor approaches 0.5 and the ratio of the current to past squared gradients. The following paragraphs explain the motivations for this adjustment.
@@ 2.3 Acceleration factor $\xi$
The adaptive coefficient has generally been influenced by both first- and second-order moments. However, alternative methods have been proposed to address challenges such as the risk of convergence to local optima and generalization issues. For instance, the Tiger(Su, 2023) and Lion(Chen et al., 2024) optimizers rely on the sign of the gradient for adaptation, while Adabelief(Zhuang et al., 2020) replaces the squared gradient with the difference between the gradient and momentum. In this study, we tackle similar issues by introducing an additional parameter, $\xi$, which is estimated using the following equation.
@@ 3.1.2 Fine-tuning
Figure 2: Testing set accuracy profiles of various optimizers on EfficientNetB0 with (a) CIFAR-10 and (b) CIFAR-100. (c) Testing set accuracy according to architectures (EfficientNetB0-6).
@@ 3.2.2 Natural language generation
The instruction fine-tuning of the LLaMA-7B(Touvron et al., 2023) using Parameter-Efficient Fine-Tuning (PEFT)(Mangrulkar et al., 2022) will serve as a benchmark test. LoRA(Hu et al., 2021) is integrated by adding branches with rank dimension 8 to each layer’s q_proj and v_proj modules. The Alpaca dataset(Taori et al., 2023), generated via self-instruct(Wang et al., 2022) from text-davinci-003, used while 2,000 samples reserved for the test. The model is trained with a batch size of 120 for 3 epochs with a cosine annealing schedule, and the performance is evaluated using maximum token lengths of 256 and 512. As shown in Fig.4, Sophia achieves lower training loss than Frankenstein but worse test loss, indicating overfitting.
@@ 4 Analysis
In this study, we present the results of CKA analysis in Fig. 9(a), comparing the performance of RMSprop with our proposed algorithm. The results clearly show that our method effectively prevents low-level features from leaking into layers near the output. Addit in fine-tuned models, a cross-optimizer analysis is shown in Fig. 9(b), where CKA maps from each epoch are compared to the final converged result. Matrix similarity is measured using the Structural Similarity Index (SSIM) metric. Notably, our algorithm consistently achieves the highest similarity values throughout the training process, demonstrating its ability to quickly form representations similar to the final converged result.
@@ 5 Conclusion
Appendix A Hyperparameter
In our experiments, we consider the milestone optimizer, including the momentum method, adaptive learning, and state-of-the-art optimizers. On the other hand, we contemplate the prevailing tasks (image classification, natural language processing) and the work that be a tough challenge for optimizers (few-shot learning). Furthermore, we introduce techniques to estimate the convergence and corresponding performance. By visualizing the loss landscape, we can measure the optimizer behavior that serves as the landscape probe. By discovering the similarity of neural network representations, we can reveal the characteristics of optimizers during training.
Admittedly, the initial learning rate and learning schedule should be tuned carefully, which would significantly affect the training result for the non-adaptive optimizers. In contrast, the main advantage of adaptive optimizers is that they serve as a general optimizer, reducing the dependency on such tuning. Thus, we set the learning rate with $10^{-3}$ except for the performance vs. network architecture research, which performs with $2\times 10^{-4}$ learning rate due to the massive network can’t converge on particular optimizers with a standard learning rate.
Appendix B Update rule comparison of various optimizers
@@
Appendix C Ablation experiments
@@
Table 6: Ablation experiments:The ablation study was tested on the CIFAR-10 dataset, with the neural network architecture being EfficientNetB0.
@@
Full 92.13
Decouple $\beta$ with LR 92.11(-0.02)
w/o $v_{t}$ 92.08(-0.05)
Fix $\beta_{2}=$0.999 92.07(-0.06)
Adam 91.97(-0.16)
Fix $\beta=$ 0.9 91.86(-0.27)
w/o $\max(v_{t-1},x_{t})$ 91.86(-0.27)
w/o EMA of $v_{t}$ 91.5(-0.63)
@@
Appendix D Adaptive overfit test
@@
Appendix E Loss landscape for energy minimization
@@
Appendix F Optimizer setting

### v2

@@ Abstract
Gradient-based optimization drives the unprecedented performance of modern deep neural network models across diverse applications. Adaptive algorithms have accelerated neural network training due to their rapid convergence rates; however, they struggle to find “flat minima” reliably, resulting in suboptimal generalization compared to stochastic gradient descent (SGD). By revisiting various adaptive algorithms’ mechanisms, we propose the Frankenstein optimizer, which combines their advantages. The proposed Frankenstein dynamically adjusts first- and second-momentum coefficients according to the optimizer’s current state to directly maintain consistent learning dynamics and immediately reflect sudden gradient changes. Extensive experiments across several research domains such as computer vision, natural language processing, few-shot learning, and scientific simulations show that Frankenstein surpasses existing adaptive algorithms and SGD empirically regarding convergence speed and generalization performance. Furthermore, this research deepens our understanding of adaptive algorithms through centered kernel alignment analysis and loss landscape visualization during the learning process. Code is available at https://github.com/acctouhou/Frankenstein_optimizer
@@ 2.2 Second momentum $v_{t}$ update with dynamic EMA
Starting with the basic RMSProp strategy, which applies a simple Exponential Moving Average (EMA) to the squared gradients, several adaptations have been developed (Appendix C). These methods depends on a fixed $\beta_{2}$ parameter throughout the training process. In contrast, this work proposes a strategy that also relies on squared gradients as its foundation but applies an EMA coefficient that adapts based on both current and historical states. This coefficient is determined by whether the $P$ factor approaches 0.5 and the ratio of the current to past squared gradients. The following paragraphs explain the motivations for this adjustment.
@@ 2.3 Acceleration factor $\xi$
The adaptive coefficient has generally been influenced by both first- and second-order moments. However, alternative methods have been proposed to address challenges such as the risk of convergence to local optima and generalization issues. For instance, the Tiger(Su, 2023) and Lion(Chen et al., 2024) optimizers rely on the sign of the gradient for adaptation, while Adabelief(Zhuang et al., 2020) replaces the squared gradient with the difference between the gradient and momentum. In this study, we tackle similar issues by introducing an additional parameter(Dubey et al., 2019), $\xi$, which is estimated using the following equation.
@@
For a more precise characterization of the optimizer’s behavior, Table 1 documents its performance metrics and the dynamics of primary update parameters across successive training stages (Early Training, Mid‑Training, Near Convergence) and in extreme cases (Gradient Spikes; Vanishing Gradients).
Table 1: Summary of conditions: behavior of $\rho$ based on gradient square $x_{t}$, factor $P$ and $\beta_{2}$.
Condition $x_{t}$ $P$ $\rho\beta_{1}$ $\beta_{2}$ Optimizer Implication
Early Training Large $\sim 0$ Large ($\sim 0.95$) $\sim 0.5$ SGD with high momentum
Mid-Training Moderate Increasing Decreasing Increasing Adam with low $\beta_{2}$
Near Convergence Small $\sim 0.5$ $\sim 0.9$ high $\beta_{2}$ Adam with high $\beta_{2}$
Gradient Spike Very large $>0.5$ $<0.9$ very low $\beta_{2}$ SGD with low momentum
Vanishing Gradients Very Small $>0.5$ $\sim 0.9$ $\approx 1$ Adam with ultra high $\beta_{2}$
@@ 3.1.2 Fine-tuning
Figure 2: Testing set accuracy profiles of various optimizers on EfficientNetB0 with CIFAR-100 (a) and (b). (c) Testing set accuracy according to architectures (EfficientNetB0-6).
@@ 3.2.2 Natural language generation
The instruction fine-tuning of the LLaMA-7B(Touvron et al., 2023) using Parameter-Efficient Fine-Tuning (PEFT)(Mangrulkar et al., 2022) will serve as a benchmark test. LoRA(Hu et al., 2021) is integrated by adding branches with rank dimension 8 to each layer’s q_proj and v_proj modules. The Alpaca dataset(Taori et al., 2023), generated via self-instruct(Wang et al., 2022) from text-davinci-003, used while 2,000 samples reserved for the test. The model is trained with a batch size of 120 for 3 epochs with a cosine annealing schedule, and the performance is evaluated using maximum token lengths of 256 and 512. As shown in Fig.4, Sophia achieves lower training loss than Frankenstein but worse test loss, indicating overfitting. In addition, the Adam optimizer is affected by grouping sequences based on their length during training, which results in performance spikes on the test set. This behavior reflects an underlying issue of training instability.
@@ 4 Analysis
In this study, we present the results of CKA analysis in Fig. 9(a), comparing the performance of RMSprop with our proposed algorithm. The results clearly show that our method effectively prevents low-level features from leaking into layers near the output. Addit in fine-tuned models, a cross-optimizer analysis is shown in Fig. 9(b), where CKA maps from each epoch are compared to the final converged result. Matrix similarity is measured using the Structural Similarity Index (SSIM) metric. It is important to note that while task performance metrics such as loss and accuracy can be clearly quantified, they do not adequately capture whether the learned representations undergo substantial changes during fine-tuning. This aspect is crucial for understanding the model’s zero-shot capability. To investigate this, we use representations pre-trained on large-scale datasets, such as ImageNet, as a strong baseline and then apply simple fine-tuning on a downstream task, such as CIFAR-100. If the model weights remain largely unchanged while still achieving strong performance on the downstream task, this suggests that the representations are stable. This stability can be assessed by measuring the CKA similarity between the final model and the initial pre-trained model. A higher similarity indicates minimal representational drift. In our experiments, we observed that models more prone to overfitting, such as those trained with Adam or RMSProp, showed greater divergence in their final representations
@@ 5 Conclusion
This study also emphasizes the recombination of various findings, showcasing how rational integration of multiple prior improvements can enhance performance. The CKA and landscape methods serve as valuable tools for analyzing learning dynamics, providing insights into the behavior of both optimizers and models. During this process, we observed that certain popular algorithms enjoy a privileged (Elhage et al., 2023) status across different fields. Furthermore, the values of our momentum coefficients, $\beta$ and $\beta_{2}$, remain within the range of 0 to 1. As a result, we did not perform a formal convergence analysis.
Acknowledgments
This work was supported by SES AI. Computational resources were partially provided by the Texas Advanced Computing Center (TACC) and the MIT SuperCloud. Additional support was provided by the “High Entropy Materials Center,” funded through the Featured Areas Research Center Program under the Higher Education Sprout Project by the Ministry of Education (MOE), Taiwan.
Appendix A More experiment detail
Regarding the from-scratch training results on ImageNet, as shown in the figure below, we ran experiments using ResNet-18 and ResNet-50 for 120 epochs. Each experiment was repeated 5 times, and only the average values are reported in the fig 10. The hyperparameters were adopted from the recommended settings in the TIMM library(Wightman, 2019). For example, for ResNet-18, we used: --batch-size 256 (across 4 GPUs), --weight-decay 1e-2, --sched step, --lr 1e-3, --epochs 120, --min-lr 1e-5, --decay-epochs 40, --amp. No specific hyperparameter search was performed, and the batch size was limited by the GTX 2080Ti 11 GB VRAM. Additionally, we examined the training behavior of Tiny ViT(Wu et al., 2022) with various optimizers, as shown in Fig. 11.
Figure 10: Test accuracy progression during training for Resnet18 and Resnet50 Frankenstein optimizer. The curves represent multiple independent runs, illustrating training stability and convergence.
Figure 11: Validation Top-1 Accuracy of Tiny ViT 5m 224 on ImageNet. Comparing performance with AdamW and Frankenstein optimizers.
Appendix B Hyperparameter
In our experiments, we conducted a comprehensive evaluation of various optimization algorithms, ranging from classical momentum-based methods to modern adaptive and state-of-the-art optimizers. We assessed their performance across standard tasks like image classification (on CIFAR-10/100) and natural language processing (using BERT fine-tuning), as well as on challenging scenarios such as few-shot learning. To gain deeper insights into optimizer behavior, we employed techniques including the visualization of the loss landscape as a probe and the analysis of neural network representation similarity.
To address concerns regarding experimental reproducibility and provide robust results, all experiments were conducted with five independent runs. The final performance metrics presented are the average results across these repeated trials to ensure reliability.
While adaptive optimizers generally offer greater robustness to hyperparameter choices, we performed systematic tuning for all optimizers on each task to ensure fair comparisons and report results based on their best achievable performance.
For the CIFAR-10/100 and NLU tasks, we conducted a hyperparameter search to identify the optimal settings for each optimizer. Tables 6 and 7 detail the learning rate search spaces explored. Other hyper-parameters, such as weight decay, were kept consistent across optimizers where applicable, or tuned individually when necessary for optimal performance.
Table 6: Hyper-parameters searching for NLU tasks.
Learning rate Batch size
$5\times 10^{-4}$ 128
$1\times 10^{-4}$
$5\times 10^{-5}$
If using SGD, lr×$\times$10
Table 7: Hyper-parameters searching for Cifar tasks.
Learning rate Batch size
$5\times 10^{-3}$ 128
$1\times 10^{-3}$ 512
$5\times 10^{-4}$
$1\times 10^{-4}$
If using SGD, lr×$\times$10
Appendix C Update rule comparison of various optimizers
@@
Appendix D Ablation experiments
@@
Table 9: Ablation experiments:The ablation study was tested on the CIFAR-10 dataset, with the neural network architecture being EfficientNetB0. Results are the average of 5 experimental runs.
@@
Full $93.17\pm 0.17$
Fix $\beta_{2}=$0.999 $92.32\pm 0.29$(-0.85)
Decouple $\beta$ with LR $92.22\pm 0.22$(-0.95)
Adam $92.12\pm 0.17$(-1.05)
Fix $\beta=$ 0.9 $92.08\pm 0.28$(-1.09)
w/o $\max(v_{t-1},x_{t})$ $91.77\pm 0.87$(-1.4)
w/o EMA of $v_{t}$ $91.74\pm 0.42$(-1.43)
@@
Appendix E Adaptive overfit test
@@
Appendix F Loss landscape for energy minimization
@@
Appendix G Optimizer setting